OpenStudy (anonymous):

Which ordered pair is not a solution of y= 4x- 9? A. (-9, 27) B. (3, -21) C. (-6, 15) D. (5, -28)

OpenStudy (anonymous):

this is on a graph?

OpenStudy (anonymous):

no, there isnt a graph for this question

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

i think its A but not sure

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

wait for texas to reply

OpenStudy (anonymous):

okay

OpenStudy (texaschic101):

just sub in your answer choices and see which one is incorrect. I will do the first one for you.. y = 4x - 9 (-9,27)....x = -9 and y = 27 lets sub 27 = 4(-9) - 9 27 = -36 - 9 27 = - 45 --- not equal...oops...maybe I shouldn't have picked the first answer choice....thats the answer

OpenStudy (texaschic101):

I was just trying to show you how to do it.....do you understand how I got the answer ?

OpenStudy (mathmate):

Are you sure the question is not \(y= \color{red}{-}4x- 9\)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

I know know you got the answer

OpenStudy (anonymous):

*how you got the answer*

OpenStudy (texaschic101):

mathmate.. I understand that the equation is not y = -4x - 9.....the -4 is not negative...but when you sub in -9 for x, it makes a negative 36.....and - 9 + (-36) does not equal 27

OpenStudy (mathmate):

@love51342 Can you check if the question is not \(y= \color{red}{-}4x- 9\)

OpenStudy (mathmate):

@texaschic101 You're working with answer A. But that's not the only choice!

OpenStudy (texaschic101):

I see that now.....my bad...if the question asks for " not correct ". then there is more then 1 answer

OpenStudy (texaschic101):

or if the equation is wrong....makes a big difference

OpenStudy (mathmate):

@texaschic101 Not really, the question asked which ordered pair (in singular), so there is a problem with the question.

OpenStudy (texaschic101):

good catch mathmate

OpenStudy (mathmate):

Hope @love51342 comes back! Thank you!